The line through (3,4) that is parallel to the line 5x+3y=9. Write the equation of each line giving answers in standard form.
Answer by mananth(16946)

The line through (3,4) that is parallel to the line 5x+3y=9. Write the equation of each line giving answers in standard form.
...
5x+3y= 9
-5x -5x
3y=-5x+9
/3
y = -5x/3 +3
..
slope = (-5/3)
...
The line parallel will have same slope.
slope = (-5/3) point = (3,4)
..
find b
4=(-5/3)*3 +b
4=-15+b
b=19
..
y = -5x/3 + 19
..
standard form
multipl;y by 3
3y=-5x+57
5x+3y-57 =0
...

...
m.ananth@hotmail.ca